# Stochastic Two-temperature Nonequilibrium Ising model

### v1

@@ I Introduction
The Ising model also provides a natural framework for exploring out-of-equilibrium behaviour. Over the years, several nonequilibrium versions of the model have been explored in different contexts. A commonly studied setup involves coupling different regions of the system to thermal baths at different temperatures. This naturally drives currents and leads to nontrivial steady-state profiles of magnetization and energy. Such arrangements have been investigated in both one [10, 11, 12, 13, 14, 15, 16] and two dimensions [17]. Generalisations involving spatially varying couplings or temperature fields have also been investigated [18, 19]. Another direction focuses on models that break detailed balance at the microscopic level—typically by introducing spin-flip rates that depend on two or more temperatures. These systems have been examined using superstatistics [20], and comparisons across different update schemes (Metropolis, Glauber, Swendsen-Wang) suggesting that certain signatures of equilibrium criticality may persist even far from equilibrium [21]. In related studies, the interplay between finite and infinite temperature baths has been shown to induce interesting steady-state features, such as nonmonotonic variation of surface tension [22]. More recently, time-dependent temperature protocols have attracted attention, where the Ising system is alternately coupled to thermal baths at different temperatures. These studies range from exploring calorimetric response in mean-field models [23] to exploring applications such as protein sequence inference [24]. However the nonequilibrium stationary behaviour of the Ising model under such stochastic temporal driving remains mostly unexplored.
@@
The paper is organised as follows. In the next section we define the model and present a brief summary of our main results. Section III focuses on the behaviour of magnetization and energy in the stationary state. The effective temperature picture in the large $\gamma$ regime is discussed in Sec. IV. We conclude with some general remarks in Sec. V
@@ II Model and Results
where $\Delta E$ is the change in the energy due to the proposed spin flip and $\beta_{\sigma}\equiv 1/T_{\sigma}=1/(T_{c}+\sigma\delta)$ denotes the inverse temperature with $\sigma=\pm 1$. The intermittent change in the temperature is incorporated via the dichotomous process $\sigma\to-\sigma$ with rate $\gamma$. Note that, $\psi_{0}$ is taken to be a constant, independent of temperature, and characterises the inverse of the typical time-scale of spin flips in the system.
@@ III Characterization of the NESS
We initialize the system in a random configuration, where each spin $s_{i}$ is chosen to be $\pm 1$ with equal probability $1/2$. Additionally, we initialize $\sigma=\pm 1$ with equal probability $1/2$, which corresponds to a temperature $T_{\sigma}=T_{c}+\sigma\delta$. The temperature remains constant for a time $\tau$ which is drawn from an exponential distribution $P(\tau)=\gamma e^{-\gamma\tau}$, during which the spins evolve via Glauber dynamics—a randomly chosen spin flips with the rate $w(\Delta E)$ defined in Eq. (2). After time $\tau$ is elapsed, $\sigma\to-\sigma$ resulting in a temperature switch. The temperature now remains constant at $T_{-\sigma}$ for a time drawn from $P(\tau)$. This dynamics eventually leads to an NESS where we measure the relevant observables.
@@
III.1 Small $\gamma$ Regime
To understand the behaviour of the NESS in the samll $\gamma$ regime, it is useful to consider a renewal approach. Since the time evolution of the system between two consecutive temperature switches follows ordinary Ising dynamics at a constant temperature, we can write a last renewal equation,
@@ III.1 Small $\gamma$ Regime
In general, it is hard to solve the integral Eq. (6) to obtain $P_{\gamma}(C,t)$ explicitly. However, in the small $\gamma$ limit, this equation can be used to obtain the stationary value of the average of any observable $\langle O\rangle$ as discussed below. For ordinary 2D Ising model with Glauber dynamics, the relaxation time $\tau_{r}\sim\psi_{0}^{-1}L^{2}$. For small values of $\gamma\ll\tau_{r}^{-1}$, the typical duration between consecutive temperature flips $\gamma^{-1}$ is much larger compared to the typical relaxation time of ordinary Ising model. Consequently, the system approaches close to equilibrium at temperatures $T_{+}$ and $T_{-}$ between successive temperature switches. In this regime, $\mathbb{P}_{T_{\sigma}}(C,t^{\prime}|C^{\prime})$ becomes independent of the initial spin configuration $C^{\prime}$ and Eq. (6) reduces to,
@@
where $\tau_{\sigma}$ denotes the relaxation time-scale at temperature $T_{\sigma}$, and $b_{\sigma}$ is an empirical constant. Substituting Eq. (9) in Eq. (8) and performing the integration over $t^{\prime}$, we get,
@@
for ordinary Ising model at $T_{\sigma}$ and extracting the parameters $\tau_{\sigma}$ and $b_{\sigma}$, which depend on the observable, by fitting it to the form given in Eq. (9). We estimate these parameters for a given $\delta$, for both energy and magnetization, from their equilibrium relaxation dynamics and compute their $\gamma$-dependence. Figure  6 compares these predicted behaviours with $\langle M\rangle_{\gamma}$ and $\langle E\rangle_{\gamma}$ obtained from numerical simulations, which show an excellent match. This validates our prediction Eq. (10) and the assumptions leading to it.
Figure 6: Small $\gamma$ regime: (a) Average magnetisation $\langle M\rangle$ and (b) average energy $\langle E\rangle$ as functions of $\gamma$ in the regime $\gamma\ll\psi_{0}$. The symbols indicate the data obtained from numerical simulations with the shaded region indicating the error margin. The solid lines correspond to Eq. (10) with the parameters $b_{\pm}$ and $\tau_{\pm}$ extracted by fitting $g_{O}^{\sigma}(t)$ to the empirical form Eq. (9) [see the insets]. Here we have used $L=32,\psi_{0}=0.01$ and $\delta=1.0$.
This analysis also explains the non-monotonic behaviour of $\langle M\rangle_{\gamma}$ and $\langle E\rangle_{\gamma}$. In the small $\gamma$ regime, at each temperature switching event $\sigma\to-\sigma$, the expectation value $\langle O(t)\rangle_{\gamma}$ starts from $\langle O\rangle_{\rm eq}^{\sigma}$, eventually reaching $\langle O\rangle_{\rm eq}^{-\sigma}$ before the next switch [see Fig. 1]. Thus, $b_{+}$ and $b_{-}$ must have opposite signs. For example, for magnetisation $b_{+}>0$ and $b_{-}<0$, while for energy $b_{+}<0$ and $b_{-}>0$. It is then easy to show that, for $\tau_{+}\neq\tau_{-}$, Eq. (10) has an extremum at some intermediate value of $\gamma$. Since, the relaxation in the ferromagnetic phase is much slower compared to that in the paramagntic phase, $\tau_{-}>\tau_{+}$, with their difference becoming larger as $\delta$ is increased. This difference in the relaxation time-scales leads to the non-monotonic variation of $\langle M\rangle_{\gamma}$ and $\langle E\rangle_{\gamma}$ in the small $\gamma$ regime, as shown in Fig. 6.
III.2 Small $\delta$ regime
@@
Since we are interested in the response coefficient of single-time observable like $M(t)$ and $E(t)$, it is useful to consider a time-antisymmetric observable of the form $O(\omega,t)=O(t)-O(0)$ in Eq. (13). In this case, the the correlations $\langle(D^{\prime})^{2}O\rangle_{\rm eq}$, $\langle(S^{\prime})^{2}O\rangle_{\rm eq}$, and $\langle D^{\prime\prime}O\rangle_{\rm eq}$ vanish due to the time-reversal invariant nature of equilibrium dynamics. Consequently, to second order in $\delta$, we have,
@@
where $H$ is the Hamiltonian of the Ising system given in Eq. (1) and $\beta_{\sigma}=1/(T_{c}+\sigma\delta)$. Since we need to compute excesses of $S(\omega)$ and $D(\omega)$ with respect to $\delta$, it suffices to consider only the $\delta$ dependent terms,
@@
Here, $\Delta E^{\sigma}(t)$ in Eq. (19) is the total energy gained by the system from the reservoir at temperature $T_{\sigma}$ during the interval $[0,t]$. Moroever, $\Delta\varepsilon_{k}(t^{\prime})$ in Eq. (20) denotes the energy cost associated with flip of the $k$-th spin in the configuration $\{C(t^{\prime}),\sigma(t^{\prime})\}$.
The excess entropy can be computed by taking derivative of Eq. (19) with respect to $\delta$ and then taking $\delta=0$.
@@
The excess frenesy can be similarly obtained from Eq. (20) and is given by,
@@
Substituting the above equations in Eq. (16), we explicitly obtain the second-order response coefficient, which can be conveniently expressed as,
@@
It is noteworthy that the net energy gained U(t)=H(t)H(0)U(t)=H(t)-H(0), where $H(t)=-\sum_{\langle ij\rangle}s_{i}(t)s_{j}(t)$ denotes the total energy of the system at time $t$. Thus, the entropic component of the susceptibility Eq. (25) can be expressed as,
@@
Using Eqs. (24)-(26) we compute the stationary response coefficients $\chi_{2,M}$ and $\chi_{2,E}$ for magnetization and energy, respectively, from numerical simulations of the equilibrium Ising model at $T_{c}$. This is compared with the response measured directly from the perturbed system $\tilde{\chi}_{2,O}=\big{(}\langle O\rangle_{\delta}-\langle O\rangle_{\rm eq}\big{)}/\delta^{2}$ in Fig. 7, which shows an excellent match with the prediction Eq. (24). From Fig. 7, it is clear that the response coefficient vanishes at an intermediate value of $\gamma$, where,
@@
It is now useful to consider what happens to the susceptibility in the limiting situations $\gamma\to 0$ and $\gamma\to\infty$. In the limit of vanishing switching rate $\gamma\to 0$, the temperature remains constant at its initial value $T_{\sigma_{0}}$, allowing us to substitute $\sigma(t^{\prime})=\sigma_{0}$ in Eq. (26), which leads to,
@@
Performing the sum over $\sigma_{0}$ and using Eq. (30) in Eq. (12), we get
@@
which infact is the small $\delta$ expansion of Eq. (10) in the $\gamma\to 0$ limit.
On the other hand, for large switching rate, $\gamma\gg\psi_{0}$, we find that the $\gamma$ dependent part of the response function $\chi_{2}^{(2)}\sim 1/\gamma$ [see Appendix A for details]. Hence, in the limit $\gamma\to\infty$, $\chi_{2}^{(2)}\to 0$ and the susceptibility becomes solely entropic in nature [see Eq. (28)]. This is suggestive of an equililibrium-like picture in the large $\gamma$ scenario, which we investigate in the following section.
IV Large $\gamma$ regime and effective temperature description
In this section, we focus on the behavior of the NESS in the large $\gamma$ regime, i.e., when $\gamma\gg\psi_{0}$. In this regime, the typical interval between two successive temperature switching events is much smaller than the typical time-scale of spin-flips. Hence, in the limit of infinite temperature switching rate, we can replace $\sigma$ by $\langle\sigma\rangle=0$ in Eq. (2) to write the effective spin-flip rate as,
@@ IV Large $\gamma$ regime and effective temperature description
Clearly, $w_{\rm eff}(\Delta E)/w_{\rm eff}(-\Delta E)=\exp(-\beta_{\infty}\Delta E)$ and Eq. (32) corresponds to a detailed balanced Glauber dynamics at an effective temperature $T_{\infty}$. Thus, in the limiting scenario of $\gamma\to\infty$, we expect the spin configuration weight to have the Boltzmann form $P(C)\sim e^{-\beta_{\infty}H(C)}$. Physically, this corresponds to a scenario where the temprature switches happens so fast that the Ising spins are not able to distinguish between the two reservoirs.
@@
Figure 8: Large $\gamma$ regime: (a) Effective temperature $T_{\text{eff}}(\gamma)$, extracted from average magnetization and energy, as functions of the flip-rate $\gamma$, for a fixed value of $\delta=0.6$. The gray shaded region indicates where the two curves coincide. The inset illustrates how $T_{\rm eff}$ is extracted from $\langle M\rangle$ vs $T$ curve for equilibrium Ising model [see Eq. (33)]. (b) Plots of $T_{\text{eff}}(\gamma)-T_{\infty}$ as functions of $\gamma$ for different values of $\delta$. The gray shaded region marks where $T_{\rm eff}(\gamma)$ is observable independent. Here we have used $L=32$ and $\psi_{0}=0.01$.
@@
Figure 9: Effective temperature $T_{\rm eff}(\gamma)$ as function of flip-rate $\gamma$, extracted from probability of spin flips corresponding to energy change $\Delta E=4$ and $\Delta E=8$. The gray dashed line indicates $T_{\rm eff}(\gamma)$ extracted from $\langle E\rangle$ vs $\gamma$ curve [see Eq. (33) and Fig. 8(a)]. The inset shows $T_{\rm eff}(\gamma)-T_{\infty}$ decays as $1/\gamma$. The gray shaded region marks where $T_{\rm eff}(\gamma)$ is observable independent. Here we have used $L=32$, $\psi_{0}=0.01$ and $\delta=0.6$.
@@
From numerical simulations, we extract $T_{\rm eff}$ for the allowed values of $\Delta E=\pm 4$ and $\pm 8$. These are compared with the $T_{\rm eff}(\gamma)$ obtained from the stationary values of the observables using Eq. (33) [see Fig. 9]. Remarkably, for $\gamma\gg\psi_{0}$, all the curves converge, indicating that the dynamics effectively does satisfy detailed balance. Thus it appears that the spin degrees, when considered independently of the bath degrees of freedom, exhibit an effective equilibrium-like behaviour in this large switching rate regime.
@@
V Conclusion

### v2

@@ 1 Introduction
The Ising model also provides a natural framework for exploring out-of-equilibrium behaviour. Over the years, several nonequilibrium versions of the model have been explored in different contexts. A commonly studied setup involves coupling different regions of the system to thermal baths at different temperatures. This naturally drives currents and leads to nontrivial steady-state profiles of magnetization and energy. Such arrangements have been investigated in both one [10, 11, 12, 13, 14, 15, 16] and two dimensions [17]. Generalisations involving spatially varying couplings or temperature fields have also been investigated [18, 19]. Another direction focuses on models that break detailed balance at the microscopic level—typically by introducing spin-flip rates that depend on two or more temperatures [20]. These systems have been examined using superstatistics [21], and comparisons across different update schemes (Metropolis, Glauber, Swendsen-Wang) suggesting that certain signatures of equilibrium criticality may persist even far from equilibrium [22]. In related studies, the interplay between finite and infinite temperature baths has been shown to induce interesting steady-state features, such as nonmonotonic variation of surface tension [23]. More recently, time-dependent temperature protocols have attracted attention, where the Ising system is alternately coupled to thermal baths at different temperatures. These studies range from exploring calorimetric response [24] and entropy production rate [25] in mean-field models to exploring applications such as protein sequence inference [26]. However, the nonequilibrium stationary behaviour of the Ising model under such stochastic temporal driving remains mostly unexplored.
@@
The paper is organised as follows. In the next section we define the model and present a brief summary of our main results. Section 3 focuses on the behaviour of magnetization and energy in the stationary state. The effective temperature picture in the large $\gamma$ regime is discussed in Sec. 4. We conclude with some general remarks in Sec. 6
@@ 2 Model and Results
where $\Delta E$ is the change in the energy due to the proposed spin flip and $\beta_{\sigma}\equiv 1/T_{\sigma}=1/(T_{c}+\sigma\delta)$ denotes the inverse temperature with $\sigma=\pm 1$. The intermittent change in the temperature is incorporated via the dichotomous process $\sigma\to-\sigma$ with rate $\gamma$. Note that, $\psi_{0}$ is taken to be a constant, independent of temperature, and characterises the inverse of the typical time-scale of spin flips in the system. Thus for a given $\sigma$ the system evolves via ordinary Glauber dynamics at a temperature $T_{\sigma}$.
@@ 3 Characterization of the NESS
We initialize the system in a random configuration, where each spin $s_{i}$ is chosen to be $\pm 1$ with equal probability $1/2$. Additionally, we initialize $\sigma=\pm 1$ with equal probability $1/2$, which corresponds to a temperature $T_{\sigma}=T_{c}+\sigma\delta$. The temperature remains constant for a time $\tau$ which is drawn from an exponential distribution $P(\tau)=\gamma e^{-\gamma\tau}$, during which the spins evolve via Glauber dynamics—a randomly chosen spin flips with the rate $w(\Delta E)$ defined in (2). After time $\tau$ is elapsed, $\sigma\to-\sigma$ resulting in a temperature switch. The temperature now remains constant at $T_{-\sigma}$ for a time drawn from $P(\tau)$. This dynamics eventually leads to an NESS where we measure the relevant observables.
@@
3.1 Small $\gamma$ Regime
To understand the behaviour of the NESS in the small $\gamma$ regime, it is useful to consider a renewal approach [27, 28]. Since the time evolution of the system between two consecutive temperature switches follows ordinary Ising dynamics at a constant temperature, we can write a last renewal equation,
@@ 3.1 Small $\gamma$ Regime
In general, it is hard to solve the integral (6) to obtain $P_{\gamma}(C,t)$ explicitly. However, in the small $\gamma$ limit, this equation can be used to obtain the stationary value of the average of any observable $\langle O\rangle$ as discussed below. For ordinary 2D Ising model with Glauber dynamics, the relaxation time $\tau_{r}\sim\psi_{0}^{-1}L^{2}$. For small values of $\gamma\ll\tau_{r}^{-1}$, the typical duration between consecutive temperature flips $\gamma^{-1}$ is much larger compared to the typical relaxation time of ordinary Ising model. Consequently, the system approaches close to equilibrium at temperatures $T_{+}$ and $T_{-}$ between successive temperature switches. In this regime, $\mathbb{P}_{T_{\sigma}}(C,t^{\prime}|C^{\prime})$ becomes independent of the initial spin configuration $C^{\prime}$ and (6) reduces to,
@@
where $\tau_{\sigma}$ denotes the relaxation time-scale at temperature $T_{\sigma}$, and $b_{\sigma}$ is an empirical constant. Substituting (9) in (8) and performing the integration over $t^{\prime}$, we get,
@@
for ordinary Ising model at $T_{\sigma}$ and extracting the parameters $\tau_{\sigma}$ and $b_{\sigma}$, which depend on the observable, by fitting it to the form given in (9). We estimate these parameters for a given $\delta$, for both energy and magnetization, from their equilibrium relaxation dynamics and compute their $\gamma$-dependence. Figure  6 compares these predicted behaviours with $\langle M\rangle_{\gamma}$ and $\langle E\rangle_{\gamma}$ obtained from numerical simulations, which show an excellent match. This validates our prediction (10) and the assumptions leading to it.
Figure 6: Small $\gamma$ regime: (a) Average magnetisation $\langle M\rangle$ and (b) average energy $\langle E\rangle$ as functions of $\gamma$ in the regime $\gamma\ll\psi_{0}$. The symbols indicate the data obtained from numerical simulations with the shaded region indicating the error margin. The solid lines correspond to (10) with the parameters $b_{\pm}$ and $\tau_{\pm}$ extracted by fitting $g_{O}^{\sigma}(t)$ to the empirical form (9) [see the insets]. Here we have used $L=32,\psi_{0}=0.01$ and $\delta=1.0$.
This analysis also explains the non-monotonic behaviour of $\langle M\rangle_{\gamma}$ and $\langle E\rangle_{\gamma}$. In the small $\gamma$ regime, at each temperature switching event $\sigma\to-\sigma$, the expectation value $\langle O(t)\rangle_{\gamma}$ starts from $\langle O\rangle_{\rm eq}^{\sigma}$, eventually reaching $\langle O\rangle_{\rm eq}^{-\sigma}$ before the next switch [see Fig. 1]. Thus, $b_{+}$ and $b_{-}$ must have opposite signs. For example, for magnetisation $b_{+}>0$ and $b_{-}<0$, while for energy $b_{+}<0$ and $b_{-}>0$. It is then easy to show that, for $\tau_{+}\neq\tau_{-}$, (10) has an extremum at some intermediate value of $\gamma$. Since, the relaxation in the ferromagnetic phase is much slower compared to that in the paramagntic phase, $\tau_{-}>\tau_{+}$, with their difference becoming larger as $\delta$ is increased. This difference in the relaxation time-scales leads to the non-monotonic variation of $\langle M\rangle_{\gamma}$ and $\langle E\rangle_{\gamma}$ in the small $\gamma$ regime, as shown in Fig. 6.
3.2 Small $\delta$ regime
@@
Since we are interested in the response coefficient of single-time observable like $M(t)$ and $E(t)$, it is useful to consider a time-antisymmetric observable of the form $O(\omega,t)=O(t)-O(0)$ in (13). In this case, the the correlations $\langle(D^{\prime})^{2}O\rangle_{\rm eq}$, $\langle(S^{\prime})^{2}O\rangle_{\rm eq}$, and $\langle D^{\prime\prime}O\rangle_{\rm eq}$ vanish due to the time-reversal invariant nature of equilibrium dynamics. Consequently, to second order in $\delta$, we have,
@@
where $H$ is the Hamiltonian of the Ising system given in (1) and $\beta_{\sigma}=1/(T_{c}+\sigma\delta)$. Since we need to compute excesses of $S(\omega)$ and $D(\omega)$ with respect to $\delta$, it suffices to consider only the $\delta$ dependent terms,
@@
Here, $\Delta E^{\sigma}(t)$ in (19) is the total energy gained by the system from the reservoir at temperature $T_{\sigma}$ during the interval $[0,t]$. Moroever, $\Delta\varepsilon_{k}(t^{\prime})$ in (20) denotes the energy cost associated with flip of the $k$-th spin in the configuration $\{C(t^{\prime}),\sigma(t^{\prime})\}$.
The excess entropy can be computed by taking derivative of (19) with respect to $\delta$ and then taking $\delta=0$.
@@
The excess frenesy can be similarly obtained from (20) and is given by,
@@
Substituting the above equations in (16), we explicitly obtain the second-order response coefficient, which can be conveniently expressed as,
@@
It is noteworthy that the net energy gained U(t)=H(t)H(0)U(t)=H(t)-H(0), where $H(t)=-\sum_{\langle ij\rangle}s_{i}(t)s_{j}(t)$ denotes the total energy of the system at time $t$. Thus, the entropic component of the susceptibility (25) can be expressed as,
@@
Using Eqs. (24)-(26) we compute the stationary response coefficients $\chi_{2,M}$ and $\chi_{2,E}$ for magnetization and energy, respectively, from numerical simulations of the equilibrium Ising model at $T_{c}$. This is compared with the response measured directly from the perturbed system $\tilde{\chi}_{2,O}=\big(\langle O\rangle_{\delta}-\langle O\rangle_{\rm eq}\big)/\delta^{2}$ in Fig. 7, which shows an excellent match with the prediction (24). From Fig. 7, it is clear that the response coefficient vanishes at an intermediate value of $\gamma$, where,
@@
It is now useful to consider what happens to the susceptibility in the limiting situations $\gamma\to 0$ and $\gamma\to\infty$. In the limit of vanishing switching rate $\gamma\to 0$, the temperature remains constant at its initial value $T_{\sigma_{0}}$, allowing us to substitute $\sigma(t^{\prime})=\sigma_{0}$ in (26), which leads to,
@@
Performing the sum over $\sigma_{0}$ and using (30) in (12), we get
@@
which infact is the small $\delta$ expansion of (10) in the $\gamma\to 0$ limit.
On the other hand, for large switching rate, $\gamma\gg\psi_{0}$, we find that the $\gamma$ dependent part of the response function $\chi_{2}^{(2)}\sim 1/\gamma$ [see Appendix A for details]. Hence, in the limit $\gamma\to\infty$, $\chi_{2}^{(2)}\to 0$ and the susceptibility becomes solely entropic in nature [see (28)]. This is suggestive of an equililibrium-like picture in the large $\gamma$ scenario, which we investigate in the following section.
4 Large $\gamma$ regime and effective temperature description
In this section, we focus on the behavior of the NESS in the large $\gamma$ regime, i.e., when $\gamma\gg\psi_{0}$. In this regime, the typical interval between two successive temperature switching events is much smaller than the typical time-scale of spin-flips. Hence, in the limit of infinite temperature switching rate, we can replace $\sigma$ by $\langle\sigma\rangle=0$ in (2) to write the effective spin-flip rate as,
@@ 4 Large $\gamma$ regime and effective temperature description
Clearly, $w_{\rm eff}(\Delta E)/w_{\rm eff}(-\Delta E)=\exp(-\beta_{\infty}\Delta E)$ and (32) corresponds to a detailed balanced Glauber dynamics at an effective temperature $T_{\infty}$. Thus, in the limiting scenario of $\gamma\to\infty$, we expect the spin configuration weight to have the Boltzmann form $P(C)\sim e^{-\beta_{\infty}H(C)}$. Physically, this corresponds to a scenario where the temprature switches happens so fast that the Ising spins are not able to distinguish between the two reservoirs.
@@
Figure 8: Large $\gamma$ regime: (a) Effective temperature $T_{\text{eff}}(\gamma)$, extracted from average magnetization and energy, as functions of the flip-rate $\gamma$, for a fixed value of $\delta=0.6$. The gray shaded region indicates where the two curves coincide. The inset illustrates how $T_{\rm eff}$ is extracted from $\langle M\rangle$ vs $T$ curve for equilibrium Ising model [see (33)]. (b) Plots of $T_{\text{eff}}(\gamma)-T_{\infty}$ as functions of $\gamma$ for different values of $\delta$. The gray shaded region marks where $T_{\rm eff}(\gamma)$ is observable independent. Here we have used $L=32$ and $\psi_{0}=0.01$.
@@
Figure 9: Effective temperature $T_{\rm eff}(\gamma)$ as function of flip-rate $\gamma$, extracted from probability of spin flips corresponding to energy change $\Delta E=4$ and $\Delta E=8$. The gray dashed line indicates $T_{\rm eff}(\gamma)$ extracted from $\langle E\rangle$ vs $\gamma$ curve [see (33) and Fig. 8(a)]. The inset shows $T_{\rm eff}(\gamma)-T_{\infty}$ decays as $1/\gamma$. The gray shaded region marks where $T_{\rm eff}(\gamma)$ is observable independent. Here we have used $L=32$, $\psi_{0}=0.01$ and $\delta=0.6$.
@@
From numerical simulations, we extract $T_{\rm eff}$ for the allowed values of $\Delta E=\pm 4$ and $\pm 8$. These are compared with the $T_{\rm eff}(\gamma)$ obtained from the stationary values of the observables using (33) [see Fig. 9]. Remarkably, for $\gamma\gg\psi_{0}$, all the curves converge, indicating that the dynamics effectively does satisfy detailed balance. Thus it appears that the spin degrees, when considered independently of the bath degrees of freedom, exhibit an effective equilibrium-like behaviour in this large switching rate regime.
@@
5 Finite-size Effects
In the previous sections, we have analysed the behaviour of the NESS using various theoretical techniques and limiting scenarios that are independent of the system size. However, so far we have relied on a fixed system size to validate these results. It is therefore natural to examine how the observables considered depend on the system size. To this end, we investigate the system-size dependence of the average magnetization and the energy in the stationary state. Figure 11 shows plots of $\langle M\rangle$ and $\langle E\rangle$ as functions of $\gamma$ for different values of system size $L$. Clearly, the non-monotonic nature of both the average magnetization and energy remain robust across system size variation. In fact, the qualitative behaviour of both the observables remain same as the system size is varied, although the magnetization shows a pronounced quantitative dependence on $L$.
The inset in Figure 11(a) shows the behaviour of $\langle M\rangle$ as a function of the system size $L$ for different values of $\gamma$. Clearly, for small $\gamma\ll\psi_{0}$, the magnetization decreases with increasing $L$, while for large $\gamma$, it becomes largely independent of $L$. This can be understood as follows. As already discussed, in the $\gamma\to 0$ limit [see (10)], the magnetization is given by, $\langle M\rangle_{\gamma}=[\langle M\rangle_{\text{eq}}^{+}+\langle M\rangle_{\text{eq}}^{-}]/2$, where $\langle M\rangle_{\text{eq}}^{\sigma}$ denotes average magnetization of equilibrium Ising model at temperature $T_{\sigma}$. In the limit of thermodynamically large system size, $\langle M\rangle_{\text{eq}}^{+}\to 0$ while $\langle M\rangle_{\text{eq}}^{-}$ attains a finite value, leading to $\langle M\rangle_{\gamma}\simeq\langle M\rangle_{\text{eq}}^{-}/2$. On the other hand, as discussed in Sec 4, for $\gamma\to\infty$, the spin-configurations in the NESS has an effective equilibrium description at a temperature $T_{\text{eff}}<T_{c}$. This implies that for large $\gamma$, the system attains an ordered state, yielding a large value of the magnetization. The average energy shows an opposite trend—for small $\gamma$, $\langle E\rangle$ increases as $L$ is increased. Moreover, the overall system size dependence of the average energy is less pronounced, as can be seen from Fig. 11(b).
Figure 11: Finite size effect: (a) Average magnetization $\langle M\rangle$ and (b) average energy $\langle E\rangle$ as functions of flip-rate $\gamma$ for different system sizes $L$. The inset in panel (a) shows plots of $\langle M\rangle$ versus $L$ for different values of $\gamma$. Here we have used $\psi_{0}=0.01$ and $\delta=1.0$.
It should be noted that both $\langle M\rangle$ and $\langle E\rangle$ exhibit little dependence on the system size in the large-$\gamma$ regime, implying that the effective temperature at large $\gamma$ is essentially independent of the system size.
6 Conclusion